3. The speed of sound in air is 1100 feet per second. The speed of sound in steel is 16500 per second. Robin, one ear pressed against the railroad track, hears a sound through the rail six seconds before hearing the same sound through the air. To the nearest foot, how far away is the source of that sound?
Answer by stanbon(75887)

The speed of sound in air is 1100 feet per second. The speed of sound in steel is 16500 per second. Robin, one ear pressed against the railroad track, hears a sound through the rail six seconds before hearing the same sound through the air. To the nearest foot, how far away is the source of that sound?
---------------
Time difference = 6 seconds
-----------------------
Rail DATA:
Distance = x ft ; Rate = 16,500 ft/sec ; Time = d/r = x/16500 sec
----------------
Air DATA:
Distance = x ft ; Rate = 11000 ft/sec ; Time = d/r = x/11000 sec
-----------------
EQUATION:
air time - rail time = 6 seconds
x/11000 - x/16500 = 6
(16500x - 11000x) = 6(11000*16500)
5500x = 6*11000*16500
x = 198,000 ft.(distance the train was when it initiated the sound heard
by Robin)
================
Cheers,
Stan H.